Maple syrup Urine disease (autosomal recessive) occurs in 1 in 380 live births in the old order Mennonite population. what is the carrier frequency of the MSUD disease allele in this population? what about in the general population, where 1 MSUD case occurs in 180,00 live births
Answer #1

Estimation of carried frequency

In the incidence of autosomal recessive disorder like MSUD calculate carrier frequency by use of some simple algebra.

q2=1/380

=1/19

p+q=1

P= 18/19

Carrier frequency can be calculated as 2*p*q

=2*18/19*1/19=~1 in 10

In case of general population

q2=1/18000

q=1/134

p+q=1

P= 133/134

Carrier frequency can calculated as 2*p*q

=2*133/134*1/134=~1 in 67